OpenStudy (anonymous):

Find the rate of change of the function h(x) = 2^x on the interval 2 ≤ x ≤ 4. The rate of change is?

OpenStudy (jdoe0001):

\(\bf h(x) = 2^x\qquad 2\le x \le 4 \\ \quad \\ \textit{"average" rate of change}\to \cfrac{h(4)-h(2)}{4-2}\)